Grade 12Physical Chemistry

600 ML OF A MIXTURE OF OZONE AND OXYGEN Weighs 1GRAM AT NTP.Calculate THE Volume OF Ozone IN THE MIXTURE

TANMOY ROY CHOUDHURY
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
 

Assume xg of Ozone . O2= 1-x g

(Total moles of gas)/22.4 = 0.6 L

{x/48 + (1-x)/32} / 22.4 = 0.6

Solve for x...